# GLTW: Joint Improved Graph Transformer and LLM via Three-Word Language for Knowledge Graph Completion

### v2

@@ Abstract
GLTW: Joint Improved Graph Transformer and LLM via Three-Word Language for Knowledge Graph Completion
Kangyang Luo, Yuzhuo Bai, Cheng Gao, Shuzheng Si, Yingli Shen Zhu Liu, Zhitong Wang, Cunliang Kong, Wenhao Li, Yufei Huang Ye Tian, Xuantang Xiong, Lei Han, Maosong Sun Tsinghua University   Tencent Robotics X
@@ 1 Introduction
Yet, they either use traditional embedding-based KGC methods Bordes et al. (2013); Lin et al. (2015); Sun et al. (2019); Balažević et al. (2019) that only consider internal links of triplets or rely on Graph Neural Networks (GNNs) Bronstein et al. (2021); Corso et al. (2020) that merely encode local subgraphs, thus both missing out on global structural knowledge. Also, LLMs, typically used for generative tasks, have long been troubled by hallucination Ji et al. (2023); Rawte et al. (2023). In contrast, the prediction targets of KGC are generally confined to the given KG, making it unwise to directly integrate LLMs into KGC tasks [^1] Notably, see Appendix A for more related works.. In short, how to encode both local and global structural information of KGs and combine it with knowledge-rich LLMs to achieve deterministic KGC remains underexplored.
To this end, we propose a novel method (named GLTW), which effectively encodes KG subgraphs with both local and global structural information and integrates LLMs in a deterministic fashion to improve the performance of KGC. Concretely, we first treat entities and relations within KG as inseparable units, adding them as tokens to the original Tokenizer, while referring to triplets as three-word sentences Guo et al. (2024). Subsequently, for each target triple, we extract a subgraph that encompasses both local and global structural information from the given training KG data (Section 3.1). To effectively process the subgraph, we introduce an improved Graph Transformer (iGT), which takes the entity and relation embeddings (initialized by a pooling operation), the relative distance matrix, and the relative distinction matrix of the subgraph as inputs, and encodes them using the enhanced graph attention mechanism (Section 3.2). Furthermore, we construct multiple positive and negative triplet samples from the subgraph, which are used to build the subgraph-based multi-classification training objective with all entities within the KG as classification objects (Section 3.3). Finally, we merge iGT with an LLM that takes KG language prompt as input (Section 3.4). To sum up, we highlight our contributions as follows:
@@ 4.2 Results Comparison
We compare the proposed methods with various KGC baselines on FB15k-237, WN18RR, and Wikidata5M, with the results shown in Table 1. The results indicate that: 1) GLTW7b consistently outperforms all competitors across all metrics, achieving overall gains of $8.5\%$ in MRR, $6.9\%$ in Hits@1, $10.2\%$ in Hits@3, and $6.1\%$ in Hits@10 compared to the second-best results (mostly from GLTW3b). Meanwhile, GLTW’s performance improves as the LLM size increases. These results demonstrate that GLTW effectively captures the characteristics of entities and relations in KGs and leverages the rich knowledge in LLMs to enhance prediction accuracy. 2) GLTW3b beats Llama-2-7b-based baseline MKGL (the most comparable method) on all metrics for FB15k-237 and WN18RR, with further improvements achieved by GLTW7b. We attribute GLTW’s advantage to its effective encoding of both local and global structural information of KGs, tailoring a suitable objective function for training subgraphs, and enabling LLMs to perceive structural information and effectively participate in entity prediction. 3) The proposed iGT consistently outstrips other GT/GNN-based baselines on FB15k-237 and WN18RR, while $g$GLM uniformly lags behind others. A detailed analysis is provided in the Ablation Study (see Section 4.3).
@@ Appendix A Related Work
KGC with LLMs. LLMs are deemed highly promising in the realm of KGC and have garnered extensive attention Ren et al. (2024); Yu et al. (2024); Pan et al. (2024); si2024selecting; Luo et al. (2024); Si et al. (2025). For instance, Yao et al. (2023); Zhu et al. (2024); Wei et al. (2024); Li et al. (2024a); Xu et al. (2024) directly perform KGC via ICL or enhance textual information in KGs to improve text-based methods. However, these methods overlook the inherent structural information of KGs, leaving LLMs unable to perceive structural knowledge. To tackle this, Zhang et al. (2024); Liu et al. (2024b); Yang et al. (2024b) integrate structural information with LLMs to boost KGC performance. Recently, MKGL Guo et al. (2024) enables LLMs to proficiently grasp entities and relations of KGs through three-word language, but how to make LLMs perceive graph information and improve the link prediction task remains an open problem. Going beyond the aforementioned methods, there are a handful of recent studies Li et al. (2024b); Xue et al. (2024); Jiang et al. (2024) on leveraging LLMs for KGC.
@@ Appendix B Complete Experimental Settings
During training, we assign distinct training schedules to different modules to fully capture the knowledge in the KG datasets. These modules include iGT Encoder, LLM, Adapter and Classification Layer. Notably, we may also train the pooling operators. When training resources are limited, we follow Guo et al. (2024) by drawing on LoRA technology Hu et al. (2021) to mitigate memory consumption. For ease of description, we divide the training modules of GLTW into three parts: iGT Encoder, LLM, and the remaining modules (referred to as "Other Modules"). Specifically, for FB15k-237 and WN18RR, we set the number of training epochs to $10$ and the gradient accumulation steps to $4$. For Wikidata5M, we set the number of training epochs to $2$ and the gradient accumulation steps to $10$. In all experiments, we used a linear learning rate schedule and the AdamW optimizer. For iGT Encoder, LLM, and Other Modules, we set the learning rates to $0.0001$, $0.00001$, and $0.001$, respectively, with warm-up rates (i.e., the proportion of warm-up steps to total training steps) of $0.02$, $0.04$, and $0.01$. Given that we used three different-sized LLMs, during training, we set the batch size per device to 16 for GLTW7b, 32 for GLTW3b, and 64 for GLTW1b over WN18RR and Wikidata5M. For FB15k-237, the batch sizes are set to 32 for GLTW7b, 64 for GLTW3b, and 128 for GLTW1b. Note that for all LLMs, we fine-tuned them using LoRA technology, with parameters set as follows: $r=32$, $dropout=0.05$, and $target\ modules=(query,value)$. We conduct our experiments on NVIDIA A800 40G GPUs with DeepSpeed+ZeRO3 and BF16. We will make the code and data publicly available upon acceptance. To ensure reliability, we report the average for each experiment over $3$ random seeds.

### v4

# GLTW: Joint Improved Graph-Transformer Encoder and LLM via Three-Word Language for Knowledge Graph Completion

@@ Abstract
GLTW: Joint Improved Graph-Transformer Encoder and LLM via Three-Word Language for Knowledge Graph Completion
Kangyang Luo, Yuzhuo Bai, Cheng Gao, Shuzheng Si, Yingli Shen Zhu Liu, Zhitong Wang, Cunliang Kong, Wenhao Li, Yufei Huang Ye Tian, Xuantang Xiong, Lei Han, Maosong Sunthanks: Corresponding author Department of Computer Science and Technology, Tsinghua University Institute for AI, Tsinghua University  Tencent Robotics X Jiangsu Collaborative Innovation Center for Language Ability
@@ 1 Introduction
Yet, they either use traditional embedding-based KGC methods Bordes et al. (2013); Lin et al. (2015); Sun et al. (2019); Balažević et al. (2019) that only consider internal links of triplets or rely on Graph Neural Networks (GNNs) Bronstein et al. (2021); Corso et al. (2020) that merely encode local subgraphs, thus missing out on global structural knowledge. Also, LLMs, typically used for generative tasks, have long been troubled by hallucination Ji et al. (2023); Rawte et al. (2023). In contrast, the prediction targets of KGC are generally confined to the given KG, making it unwise to directly integrate LLMs into KGC tasks [^1] Notably, see Appendix A for more related works.. In short, how to encode both local and global structural information of KGs and combine it with knowledge-rich LLMs to achieve deterministic KGC remains underexplored.
To this end, we propose a novel method (named GLTW), which effectively encodes KG subgraphs with both local and global structural information and integrates LLMs in a deterministic fashion to improve the performance of KGC. Concretely, we first treat entities and relations within KG as inseparable units, adding them as tokens to the original Tokenizer, while referring to triplets as three-word sentences Guo et al. (2024). Subsequently, for each target triple, we extract a subgraph that encompasses both local and global structural information from the given training KG data (Section 3.1). To effectively process the subgraph, we introduce an improved Graph Transformer (iGT), which takes the entity and relation embeddings (initialized by a pooling operation), the relative distance matrix, and the relative distinction matrix of the subgraph as inputs, and encodes them using the enhanced graph attention mechanism (Section 3.2). Furthermore, we construct multiple positive and negative triplet samples from the subgraph, which are used to build the subgraph-based multi-classification training objective with all entities within the KG as classification objects (Section 3.3). Finally, we merge iGT with LLM that takes KG language prompt as input (Section 3.4). To sum up, we highlight our contributions as follows:
@@ 4.2 Results Comparison
We compare the proposed methods with various KGC baselines on FB15k-237, WN18RR, and Wikidata5M, with the results shown in Table 1. The results indicate that: 1) GLTW7b consistently outperforms all competitors across all metrics, achieving overall gains of $8.5\%$ in MRR, $6.9\%$ in Hits@1, $10.2\%$ in Hits@3, and $6.1\%$ in Hits@10 compared to the second-best results (mostly from GLTW3b). Meanwhile, GLTW’s performance improves as the LLM size increases. These results demonstrate that GLTW effectively captures the characteristics of entities and relations in KGs and leverages the rich knowledge in LLMs to enhance prediction accuracy. 2) GLTW3b beats Llama-2-7b-based baseline MKGL (the most comparable method) on all metrics for FB15k-237 and WN18RR, with further improvements achieved by GLTW7b. We attribute GLTW’s advantage to its effective encoding of both local and global structural information of KGs, tailoring a suitable objective function for training subgraphs, and enabling LLMs to perceive structural information and effectively participate in entity prediction. 3) The proposed iGT consistently outstrips other GT/GNN-based baselines on FB15k-237 and WN18RR, while $g$GLM uniformly lags behind others. A detailed analysis is provided in the Ablation Study.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Science and Technology Major Project (2020AAA0106502), the National Natural Science Foundation of China (No.T2341003) and a grant from the Guoqiang Institute, Tsinghua University.
@@ Appendix A Related Work
KGC with LLMs. LLMs have been explored by researchers for various tasks due to their powerful emergent capabilities Luo et al. (2024); Yu et al. (2024); Tan et al. (2024); Si et al. (2025); Liu et al. (2025). Recently, LLMs are deemed highly promising in the realm of KGC and have garnered extensive attention Ren et al. (2024); Pan et al. (2024). For instance, Yao et al. (2023); Zhu et al. (2024); Wei et al. (2024); Li et al. (2024a); Xu et al. (2024) directly perform KGC via ICL or enhance textual information in KGs to improve text-based methods. However, these methods overlook the inherent structural information of KGs, leaving LLMs unable to perceive structural knowledge. To tackle this, Zhang et al. (2024); Liu et al. (2024b); Yang et al. (2024b) integrate structural information with LLMs to boost KGC performance. Recently, MKGL Guo et al. (2024) enables LLMs to proficiently grasp entities and relations of KGs through three-word language, but how to make LLMs perceive graph information and improve the link prediction task remains an open problem. Going beyond the aforementioned methods, there are a handful of recent studies Li et al. (2024b); Xue et al. (2024); Jiang et al. (2024) on leveraging LLMs for KGC.
@@ Appendix B Complete Experimental Settings
During training, we assign distinct training schedules to different modules to fully capture the knowledge in the KG datasets. These modules include iGT Encoder, LLM, Adapter and Classification Layer. Notably, we may also train the pooling operators. When training resources are limited, we follow Guo et al. (2024) by drawing on LoRA technology Hu et al. (2021) to mitigate memory consumption. For ease of description, we divide the training modules of GLTW into three parts: iGT Encoder, LLM, and the remaining modules (referred to as "Other Modules"). Specifically, for FB15k-237 and WN18RR, we set the number of training epochs to $10$ and the gradient accumulation steps to $4$. For Wikidata5M, we set the number of training epochs to $2$ and the gradient accumulation steps to $10$. In all experiments, we used a linear learning rate schedule and the AdamW optimizer. For iGT Encoder, LLM, and Other Modules, we set the learning rates to $0.0001$, $0.00001$, and $0.001$, respectively, with warm-up rates (i.e., the proportion of warm-up steps to total training steps) of $0.02$, $0.04$, and $0.01$. Given that we used three different-sized LLMs, during training, we set the batch size per device to 16 for GLTW7b, 32 for GLTW3b, and 64 for GLTW1b over WN18RR and Wikidata5M. For FB15k-237, the batch sizes are set to 32 for GLTW7b, 64 for GLTW3b, and 128 for GLTW1b. Note that for all LLMs, we fine-tuned them using LoRA technology, with parameters set as follows: $r=32$, $dropout=0.05$, and $target\ modules=(query,value)$. Note that in specific experiments, the element functions $f_{1}$ and $f_{2}$ are essentially learnable embedding layers, which encode numbers into corresponding vectors. Meanwhile, the element functions of matrix $P$ and matrix $D$ are independent of each other (except for the G2G relative position).